1/2

Christian Morgenstern, Christian Morgenstern,

Christian Morgenstern,

El Hechicero The Werewolf The Banshee (An Approach)

tradukita de Hans Eiseneck tradukita de Alexander Gross

tradukita de Max Knight

Un hechicero, con espo-

sa vino un día a la fosa de su maestro, y al pro-

fesor pidió: “¿Me cuentas?

¡Por favor!”

A Werewolf, troubled by his name, Left wife and brood one night and

came To a hidden graveyard to enlist The aid of a long-dead philologist.

One night, a banshee slunk

away from mate and child, and in the

gloom went to a village teacher’s tomb, requesting him: “Inflect me,

pray.”

El profesor sin vacilar subió y comenzó a hab-

lar. Desde allá de su letrero se dirigió al hechicero:

“Oh sage, wake up, please don’t be-

rate me,” He howled sadly, “Just conjugate

me.” The seer arose a bit unsteady Yawned twice, wheezed once, and

then was ready.

The village teacher climbed up

straight upon his grave stone with its

plate and to the apparition said who meekly knelt before the

dead:

“El hechiuno” le explica, “el hechidos” después in-

dica, “el hechitres” dice al fin, “con eso basta ¡chi-

quitín!”

“Well, ‘Werewolf’ is your plural past, While ‘Waswolf’ is singularly cast: There’s ‘Amwolf’ too, the present

tense, And ‘Iswolf,’ ‘Arewolf’ in this same

sense.”

“The banSHEE, in the subject’s

place; the banHERS, the possessive

case. The banHER, next, is what

they call objective case–and that is all.”

El hechicero, conmovido por lo que capta su oído, al maestro ruega, encan-

tado, que le adjunte un que-

brado.

“I know that–I’m no mental cripple– The future form and participle Are what I crave,” the beast replied. The scholar paused–again he tried:

The banshee marveled at the

cases and writhed with pleasure, ma-

king faces, but said: “You did not add, so

far, the plural to the singular!”

Pero el profesor confiesa que nada sabe con certe-

za: “Hay muchos hechos.

Hechiceros únicamente hay ente-

ros.”

“A ‘Will-be-wolf?’ It’s just too long: ‘Shall-be-wolf?’ ‘Has-been-wolf?’

Utterly wrong! Such words are wounds beyond all

suture– I’m sorry, but you have no future.”

The teacher, though, admitted

then that this was not within his ken. “While bans are frequent”, he

advised, “A she cannot be plurized.”

…  …  …

2/2

El brujo se volvió muy

triste: “Mi cara mitad, ¿no la

viste?” Ya que no era sabio con gratitud se despidió.

The Werewolf knew better–his sons

still slept At home, and homewards now he

crept, Happy, humble, without apology For such folly of philology.

The banshee, rising clammily, wailed: “What about my fami-

ly?” Then, being not a learned crea-

ture, said humbly “Thanks” and left

the teacher.

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en hispana de Hans Eiseneck.

Arg-825-1656 (2013-02-13 18:46:00)

Tiu ĉi trsduko en la kastilian lingvon troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html .

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en la Anglan de Alexander Gross.

Arg-825-1657 (2013-02-13 18:39:45)

Tiu ĉi angligo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html .

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-

31) en la Anglan de Max Knight (civila nomo: Max Kühnel, ∗1909-06-08 – †1993-08-31).

Arg-825-1660 (2013-02-13 22:36:06)

Tiu ĉi angligo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html . Pri la tradukinto Max Knight vidu: http: // www. jbeilharz. de/ morgenstern/ morgenstern_poems. html kaj http: // www. 123people. de/ ext/ frm? ti= ersonensuche% 20telefonbuch&search_term= m x% 20k% C3% BChnel&search_ country= DE&st= suc e% 20nach% 20personen&target_url= aHR0cDovL d3dy5saXRlcmF0dXJoYXVzLmF0L2luZGV4LnBocD9pZD02ODc4 .

http://www.poezio.net/version?poem-id=825&version-id=1656,1657,1660